A gas mixture contains 4.0 g of CO(g), 11.0 g of CO₂(g) and 2.0 g of CH4 (g) in a 50.0 - L

vessel. Calculate the partial pressure of each gas in the mixture.


1
Expert's answer
2021-12-14T15:16:04-0500

The partial pressure of each gas in a mixture is proportional to its mole fraction.


n = m / M

M (CO) = 28 g/mol

M (CO2) = 44 g/mol

M (CH4) = 16.04 g/mol


n (CO) = 4 / 28 = 0.14 mol

n (CO2) = 11 / 44 = 0.25 mol

n (CH4) = 2 / 16.04 = 0.13 mol


Sum of n = 0.14 + 0.25 + 0.13 = 0.52 moles

CO mole fraction = 0.14 / 0.52 = 0.27

CO2 mole fraction = 0.25 / 0.52 = 0.48

CH4 mole fraction = 0.13 / 0.52 = 0.25